OpenStudy (anonymous):

George is 4 times Dorothy's age. If George is 56, how old is Dorothy? Let d = Dorothy's age.

OpenStudy (anonymous):

56/4

OpenStudy (anonymous):

Let d = dorothy's age and g = George's age. g = 4d d = g/4 d = 56/ 4 d = 14 years old.